1. How high would water rise in the pipes of a building if the water pressure gauge shows the pressure at the ground floor to be 280KPa orKN/m^2.




1
Expert's answer
2021-01-26T12:41:17-0500

"p=\rho gh\to h=\frac{p}{\rho g}=\frac{280000}{1000\cdot9.8}=28.57(m)" . Answer